# ESVO2: Direct Visual-Inertial Odometry with Stereo Event Cameras

### v1

@@ VI-A Public Datasets Used
Our system is tested on five publicly available datasets with stereo event cameras and IMUs, which feature different event-camera types, spatial resolutions, and scene geometries, including DSEC [68], VECtor [69], MVSEC [70], rpg [71] and TUM-VIE [72]. DSEC is a large-scale dataset capturing outdoor driving scenes, and the others feature small indoor scenes with different sensor resolutions. Sequences in MVSEC are recorded using a stereo event camera mounted on drones, whereas a hand-held one is used in VECtor, rpg, and TUM-VIE. The characteristics of each dataset are summarized in Tab. IV.
@@
Note that some of the other released datasets using stereo event cameras are not applicable to our evaluation. For example, an unsatisfactory rectification is witnessed in the HKU dataset [2] due to inaccuracies in the extrinsic calibration. The ECMD dataset [73] exhibits a significant discrepancy in the streaming rate of events between the left and right event cameras, thus breaking the assumption of temporal coherence between the two cameras. Last but not the least, the M3ED dataset [74] employs event cameras with a spatial resolution of 1280 720 pixels. As discussed in [75], the higher the spatial resolution, the more sensitive each pixel becomes to brightness changes. In urban environments, the excessive number of events recorded in this dataset results in a lack of clear structure information in event images, especially in the presence of trees and green belts. This may result in failures of the proposed direct method.
@@ VI-B Comparison of Mapping: ESVO vs Ours
We compare our mapping results against those of the original ESVO pipeline [1] to demonstrate the performance improvement of our system in mapping. As illustrated in Fig. 5, the results of ESVO exhibit that horizontal structures in the environment are either not reconstructed or inaccurate in depth estimates. This is because the static-stereo method cannot precisely recover the depth of structures that are parallel to the stereo-camera baseline. In comparison, our system additionally incorporates a temporal stereo method, which allows for accurate depth estimation of structures parallel to the stereo-camera baseline. Besides, the sampled points from AA better capture edge structures with less noise (see last row of Fig. 5). All of these lead to better performance in terms of reconstruction completeness and local smoothness.
@@ VI-C Full System Evaluation
Dataset Sequence ESVO [1] ES-PTAM [53] ESIO [2] ESVIO [50] ICRA’24 [49] Ours rpg box monitor bin desk reader – – failed failed – – MVSEC indoor_1 indoor_2 – – – – indoor_3 – – failed failed indoor_4 failed failed – – – – failed failed DSEC city04_a 5.59 city04_b city04_c city04_d – – city04_e city04_f city11_a – – city11_b – – – – VECtor robot_normal – – – – corner_slow – – – – hdr_normal – – – – sofa_normal failed failed – – – – failed failed desk_normal – – failed failed – – TUM-VIE 1d_trans failed failed – – 3d_trans failed failed – – 6d_trans failed failed – – failed failed desk failed failed – – desk2 failed failed – –
@@
Dataset Sequence ESVO [1] ES-PTAM [53] ESIO [2] ESVIO [50] ICRA’24 [49] Ours rpg box monitor bin desk reader – – failed failed – – MVSEC indoor_1 indoor_2 – – – – indoor_3 – – failed failed indoor_4 failed failed – – – – failed failed DSEC city04_a city04_b city04_c city04_d – – city04_e city04_f city11_a – – city11_b – – – – VECtor robot_normal – – – – corner_slow – – – – hdr_normal – – – – sofa_normal failed failed – – – – failed failed desk_normal – – failed failed – – TUM-VIE 1d_trans failed failed – – 3d_trans failed failed – – 6d_trans failed failed – – failed failed desk failed failed – – desk2 failed failed – –
To quantitatively evaluate the motion estimation results, we use two standard evaluation metrics: relative pose error (RPE) and absolute trajectory error (ATE) [76]. The results of all methods can be found in Tabs. VII and VII, in which the best results are highlighted in bold. In general, the proposed method (“Ours”) outperforms the others in terms of RPE and ATE. Our method resolves all the remaining issues of ESVO [1], thus leading to more accurate trajectory estimates than ESVO.
@@
Figures 7 and 7 provide a qualitative comparison of the trajectories estimated by ESVO and our method on sequences from rpg, MVSEC, and DSEC data (To deliver a compact and clear visualization, only these trajectories are displayed). Since DSEC does not provide ground-truth trajectories, we use the result of a LiDAR-IMU–based method [77] as ground truth. In short, the trajectories obtained by our method align better with the ground truth, as evidenced by the quantitative analysis.
@@ VI-E Ablation Studies
We conduct several ablation studies to assess the benefits brought by the new functions and modules proposed in this work. Specifically, we use the DSEC dataset [68] and compare the performance of our system (denoted as “reference” in Tab. VIII) with the following features disabled, respectively.
@@ VI-F Computational Efficiency
To assess the computational efficiency, we compare the runtime of ESVO [1], ICRA’24 [49], and the proposed system using a desktop with an Intel Core i7-14700k CPU, as shown in Tab. IX. The tests are performed on the DSEC dataset. All three systems are implemented using hyper-threading technology, and the number of threads occupied by each node is declared inside the parentheses following the node name. The average runtime of each function is profiled, and the numbers in parenthesis (hyper-parameters) denote the amount of data processed by each function. It should be noted that the runtime of our system is primarily influenced by the spatial resolution of the event cameras, proportionally to which the number of points sampled from AA is set. All other hyper-parameters (e.g., the number of points used by each function) are set accordingly, aiming at best performance. As a result, the event streaming rate has minimal influence on our system’s runtime.
@@
Most of the modifications are made in the mapping module because its performance is mostly relevant to the accuracy and efficiency of the entire system. The improvement in terms of mapping efficiency is due to the following: () The number of points required by the proposed mapping module is significantly reduced compared to ESVO because the input points are no longer raw events but precise contour points extracted by the applied AA and contour-point sampling method; () The proposed fast block matching speeds up the static-stereo operation; () We get rid of the depth refinement because we find that the nonlinear optimization does not bring notable improvements in mapping accuracy when the accurate contour points are fed as input; () The newly added temporal-stereo operation causes a runtime increment (5 ms). Nevertheless, it is worthwhile as it is the key to the improvement of the overall mapping quality. Moreover, the optional regularization operation is no longer used because it is time-consuming and has little effect on the front-end’s performance. The “others” entry in Tab. IX includes general operations in the code that cannot be named as functions, such as data transfer and recycle. The subtotal runtime of the mapping module, shown in Fig. 11, demonstrates that our mapping module can stably operate in real time at 20 Hz on input stereo events of VGA resolution. Last but not the least, our back-end runs in another independent thread, taking 5 ms for each update of the optimization variables in the sliding window.

### v3

@@ VI-A Public Datasets Used
Our system is tested on five publicly available datasets with stereo event cameras and IMUs, which feature different event-camera types, spatial resolutions, and scene geometries, including DSEC [82], VECtor [83], MVSEC [84], rpg [85] and TUM-VIE [86]. DSEC is a large-scale dataset capturing outdoor driving scenes, and the others feature small indoor scenes with different sensor resolutions. Sequences in MVSEC are recorded using a stereo event camera mounted on drones, whereas a hand-held rig is used in VECtor, rpg, and TUM-VIE. The characteristics of each dataset are summarized in Tab. IV.
@@
Note that some of the other released datasets using stereo event cameras are not applicable to our evaluation. For example, an unsatisfactory rectification is witnessed in the HKU dataset [2] due to inaccuracies in the extrinsic calibration. The ECMD dataset [87] exhibits a significant discrepancy in the streaming rate of events between the left and right event cameras, thus breaking the assumption of temporal coherence between the two cameras. Last but not the least, the “Kirschbaum problem” [88] is witnessed in the M3ED dataset [89], which employs event cameras with a spatial resolution of 1280 720 pixels and collects data in urban environments filled with high-frequency texture patterns (e.g., trees and lawns). According to [90], the higher the spatial resolution, the more sensitive each pixel becomes to brightness changes. Therefore, such a sensor-scene configuration results in an excessive number of events recorded in this dataset, leading to a lack of clear structure information in the spatio-temporal profile of events. This may result in failures of the proposed direct method.
@@ VI-B Comparison of Mapping: ESVO vs Ours
We compare our mapping results against those of the original ESVO pipeline [1] to demonstrate the performance improvement of our system in mapping. As illustrated in Fig. 5, the results of ESVO exhibit that horizontal structures in the environment are either not reconstructed or inaccurate in depth estimates. This is because the static-stereo method cannot precisely recover the depth of structures that are parallel to the stereo-camera baseline. In comparison, our system additionally incorporates a temporal stereo method, which allows for accurate depth estimation of structures parallel to the stereo-camera baseline (see Sec. IV-B1). Besides, the sampled points from AA better capture edge structures with less noise (see last row of Fig. 5). All of these lead to better performance in terms of reconstruction completeness and local smoothness.
@@ VI-C Full System Evaluation
Dataset Sequence ESVO [1] ES-PTAM [66] ESIO [2] ESVIO [63] ICRA’24 [42] Ours rpg box monitor bin desk reader – – failed failed – – MVSEC indoor_1 indoor_2 – – – – indoor_3 – – failed failed indoor_4 failed failed – – – – failed failed DSEC city04_a 5.59 city04_b city04_c city04_d – – city04_e city04_f city09_b failed failed – – city11_a – – city11_b – – – – VECtor robot_normal – – – – robot_fast failed failed – – failed failed – – failed failed corner_slow – – – – hdr_normal – – – – sofa_normal failed failed – – – – failed failed desk_normal – – failed failed – – TUM-VIE 1d_trans failed failed – – 3d_trans failed failed – – 6d_trans failed failed – – failed failed desk failed failed – – desk2 failed failed – –
@@
Dataset Sequence ESVO [1] ES-PTAM [66] ESIO [2] ESVIO [63] ICRA’24 [42] Ours rpg box monitor bin desk reader – – failed failed – – MVSEC indoor_1 indoor_2 – – – – indoor_3 – – failed failed indoor_4 failed failed – – – – failed failed DSEC city04_a city04_b city04_c city04_d – – city04_e city04_f city09_b failed failed – – city11_a – – city11_b – – – – VECtor robot_normal – – – – robot_fast failed failed – – failed failed – – failed failed corner_slow – – – – hdr_normal – – – – sofa_normal failed failed – – – – failed failed desk_normal – – failed failed – – TUM-VIE 1d_trans failed failed – – 3d_trans failed failed – – 6d_trans failed failed – – failed failed desk failed failed – – desk2 failed failed – –
To quantitatively evaluate the motion estimation results, we use two standard evaluation metrics: relative pose error (RPE) and absolute trajectory error (ATE) [92]. The results of all methods can be found in Tabs. VII and VII, in which the best results are highlighted in bold. In general, the proposed method (“Ours”) outperforms the others in terms of RPE and ATE. Our method resolves most of the remaining issues of ESVO [1], thus leading to more accurate trajectory estimates than ESVO.
@@
Additionally, ESVO2 demonstrates excellent performance in both high-speed motion and HDR scenes. The vector_robot_fast sequence is captured indoors with a stereo event camera under aggressive motion, and only our method can operate on it and achieve good results. Both dsec_city_09b and vector_hdr_normal are HDR sequences of nighttime driving and small-scale indoor scenarios, respectively. Benefiting from the accurate contour points sampling strategy, ESVO2 achieves the best performance on both of these sequences.
Figures 7 and 7 provide a qualitative comparison of the trajectories estimated by ESVO and our method on sequences from rpg, MVSEC, and DSEC data (To deliver a compact and clear visualization, only these trajectories are displayed. A more complete qualitative comparison is provided in the Appendix). Since DSEC does not provide ground-truth trajectories, we use the result of a LiDAR-IMU–based method [91] as ground truth. In short, the trajectories obtained by our method align better with the ground truth, as evidenced by the quantitative analysis.
@@
| Sequence | Window size 8 | Window size 12 | w/o IMU | Reference (size 5) | ||||
| dsec_city04_a | 3.33 | 0.56 | ||||||
| dsec_city04_c | 10.26 | 5.09 | ||||||
| dsec_city04_d | 8.87 | 5.04 | ||||||
| dsec_city04_e | 3.54 | 0.76 | ||||||
| dsec_city04_f | 5.00 | 3.28 | ||||||
| dsec_city11_a | 1.85 | 0.48 | ||||||
| dsec_city11_b | 10.61 | 4.41 | ||||||
@@ VI-E Ablation Studies
We conduct several ablation studies to assess the benefits brought by the new functions and modules proposed in this work (Secs. VI-E1 to VI-E3). Additionally, we investigative the sensitivity with respect to the size parameter in the sliding window and with respect to running our pipeline with/without the IMU input (Sec. VI-E4). In all studies, we use the DSEC dataset [82] and compare the performance of our system (denoted as “reference” in Tabs. VIII and IX) with the following features disabled, respectively.
@@ VI-E3 Effect of the Nonlinear Refinement of Depth
VI-E4 Sensitivity with respect to the Size of the Sliding Window and Necessity of using an IMU
To investigate how the performance changes when using different sizes of the sliding window and running the pipeline without any IMU measurements, we conduct another ablation study, as reported in Tab. IX. The “reference” group denotes the results obtained under the default configuration, namely choosing a window size of 5 and using the IMU measurements. Compared to the “reference” group, the pose-estimation accuracy oscillates slightly up and down when using a window size of 8 or 12. This indicates that the choice of using different window sizes does not have a significant effect on the results. Our empirical selected window size aims to strike a balance between efficiency and accuracy. Second, it is clearly seen that the accuracy of the pose estimation drops notably when the IMU is not used, demonstrating the importance of introducing the IMU into the sensor suite.
@@ VI-F Computational Efficiency
To assess the computational efficiency, we compare the runtime of ESVO [1], ICRA’24 [42], and the proposed system using a desktop with an Intel Core i7-14700k CPU, as shown in Tab. X. The tests are performed on the DSEC dataset. All three systems are implemented using hyper-threading technology, and the number of threads occupied by each node is declared inside the parentheses following the node name. The average runtime of each function is profiled, and the numbers in parenthesis (hyper-parameters) denote the amount of data processed by each function. It should be noted that the runtime of our system is primarily influenced by the spatial resolution of the event cameras, proportionally to which the number of points sampled from AA is set. All other hyper-parameters (e.g., the number of points used by each function) are set accordingly, aiming at best performance. As a result, the event streaming rate has minimal influence on our system’s runtime.
@@
Most of the modifications are made in the mapping module because its performance is mostly relevant to the accuracy and efficiency of the entire system. The improvement in terms of mapping efficiency is due to the following: () The number of points required by the proposed mapping module is significantly reduced compared to ESVO because the input points are no longer raw events but precise contour points extracted by the applied AA and contour-point sampling method; () The proposed fast block matching speeds up the static-stereo operation; () We get rid of the depth refinement because we find that the nonlinear optimization does not bring notable improvements in mapping accuracy when the accurate contour points are fed as input; () The newly added temporal-stereo operation causes a runtime increment (5 ms). Nevertheless, it is worthwhile as it is the key to the improvement of the overall mapping quality. Moreover, the optional regularization operation is no longer used because it is time-consuming and has little effect on the front-end’s performance. The “others” entry in Tab. X includes general operations in the code that cannot be named as functions, such as data transfer and recycle. The runtime of the mapping module demonstrates that our mapping module can stably operate in real time at 20 Hz on input stereo events of VGA resolution. Last but not the least, our back-end runs in another independent thread, taking 5 ms for each update of the optimization variables in the sliding window.
Considering that the mapping operation typically consumes the biggest portion of the computing resources in an event-based VO pipeline, we compare the mapping module’s runtime performance of four open-source direct methods. To do this, we record the mapping module’s runtime of every method on all VGA-resolution datasets and visualize their statistics using a box plot (Fig. 11). It is clearly seen that only the median runtime of our method is below the -ms bar, ensuring real-time performance on data of VGA resolution. Additionally, we assess the runtime of a feature-based method (ESIO [2]) using the same hardware. Using the recommended configuration by the authors, the average cost of the most time-consuming back-end module in ESIO is 37.34 ms. This indicates that the proposed direct method is on par with state-of-the-art feature-based methods in terms of computational efficiency; meanwhile, as shown in Secs. VI-C and VI-D, our method outperforms others in accuracy.
@@ VI-G Limitations
Besides, the employed image-like representations have limitations on further (fully) exploiting the high-speed property of event cameras. We notice the existence of several alternative representations of event data [93]. For example, an early work interprets an event stream as 3D point clouds [94] in the spatio-temporal domain, from which normal flow can be computed by fitting local planes. The computation is asynchronous and efficient (closed-form solution), and the resulting normal flow can be used for state estimation [95], leading to ultra-frame-rate update rate. However, this map-free design [95] is an open-loop scheme, outperforming in speed by sacrificing accuracy. As for other representations, such as the spatio-temporal graph of events, they are typically discussed in the task of object recognition [93] but not in the context of event-based SLAM.